OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1120. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many two-step equations auditioned today? a. 1300 b.1120 c.708 d.412

OpenStudy (diyadiya):

Let the number of 1 step equations be \(x\) and 2 step equations be \(y\)

OpenStudy (anonymous):

ok: # of one step equations -- X # of two step equations -- Y X + Y = 1150 (the total eliminated) 3X - 2Y = 1300 (second equation) ok now find the value of Y in terms of X from the first equation EX. if 4X + 2Y = 600, the value of Y in terms of X is 2Y = 600 - 4X ----> Y = 300 - 2X Now substitue the new value of Y into the second equation and solve for X